OpenStudy (anonymous):

empirical formula of a compound containing 1.732 g carbon, 0.289 g Hydrogen, and 0.45 g oxygen

OpenStudy (jebonna):

To work out empirical formula, the first thing you have to do is work out the moles of each element in your question. You can do this by using this equation: moles = mass/molar mass So, to find moles of carbon, you'd do: 1.732g/12g mol-1 = 0.144mol (3 significant figures) Do this for the rest of the elements in your question (hydrogen and oxygen). Hydrogen: 0.289g/1g mol-1 = 0.289mol (3 significant figures) Oxygen: 0.45g/16g mol-1 = 0.0281mol (3 significant figures) Now, you have to divide the smallest moles by every value of moles you have worked out. So first, identify which is the smallest number of moles. That is 0.0281mol. So, now we divide 0.0281 by every value of moles we have worked out: Carbon: 0.144mol/0.0281mol = 5 (rounded down) Hydrogen: 0.289mol/0.0281mol = 10 (rounded down) Oxygen: 0.0281mol/0.0281mol = 1 Those numbers you have worked out are now how many of each element are in the chemical formula. This is your empirical formula as it is the lowest possible ratio of a compound possible. So your answer is: \[C _{_{5}}H _{10}O\] I hope this helps! :)